# Physics-Inspired Modeling and Content Adaptive Routing in an Infrared Gas Leak Detection Network

### v1

@@ 4.5.2 Effective Receptive Field (ERF) Analysis Of Physics–Edge Hybrid Backbone
To further verify the physical interpretability and feature propagation capability of the proposed Physics–Edge Hybrid Backbone, we conducted an effective receptive field (ERF) visualization experiment to compare it with a conventional ResNet-based backbone under identical training and inference settings, as shown in Figure. 7. Specifically, we employed the gradient-based ERF analysis method, where the gradient of the central activation with respect to the input pixels was accumulated and visualized as a heatmap (Ding et al., 2022). The visualization highlights high-response regions that contribute most significantly to the central neuron activation, thus reflecting the spatial extent and distribution of the model’s receptive field.
@@
Based on the quantitative statistics of the effective receptive field (ERF) distributions, a clear difference can be observed between the proposed Physics-Edge Hybrid Backbone and the conventional ResNet backbone. As shown in Figure. 8, both the bar and line plots illustrate the relationship between the high-contribution area ratio and the activation threshold. Specifically, under low thresholds (e.g., 0.2 and 0.3), Physics-Edge Hybrid Backbone already achieves slightly higher area ratios than ResNet, indicating a broader response range even in the early activation stage. As the threshold increases to 0.5 and 0.99, this gap becomes more pronounced — the high-contribution area ratio of Gas Block reaches 0.047 and 0.340, compared with 0.036 and 0.255 for ResNet, respectively. This trend reveals that the proposed Gas Block possesses a larger and more spatially continuous receptive field, allowing information to propagate more effectively across distant regions. Consequently, the network is better able to integrate global contextual cues while preserving fine structural details, consistent with the convection–diffusion-Inspired design of Physics-Edge Hybrid Backbone.
@@ 4.5.3 Backbone-Level Comparison With Representative Architectures
To comprehensively evaluate the effectiveness of the proposed Physics-Edge Hybrid Backbone, we conducted a comparative study against several representative models, including EfficientNet (Tan and Le, 2019), MambaOut (Yu and Wang, 2025), RepViT (Wang et al., 2024), CSWin Transformer (Dong et al., 2022),ConvNeXtV2 (Woo et al., 2023), RMT (Fan et al., 2024), EfficientFormerV2 (Li et al., 2023), LSKNet (Li et al., 2023), and VanillaNet (Chen et al., 2023), highlighting the performance improvements achieved by our method, as shown in Figure. 9.
@@ 4.5.4 Visualization Analysis Of AGPEO Edge Representation
To validate the feature extraction capability of the AGPEO module, we present a comparative visualization of the original infrared image, the gradient feature map, the phase feature map, and the fused result, as shown in Figure. 10. It can be observed that the gradient branch, employing multi-directional Sobel operators, effectively highlights prominent edge responses, emphasizing both the gas leak regions and the structural contours of the equipment. However, due to the presence of temperature gradients and noise in the infrared image background, the gradient feature map also exhibits noticeable spurious edges in some background areas, resulting in partially fragmented boundaries. The phase branch utilizes Gabor convolutional kernels to perform local phase modulation, producing phase feature maps that are smoother in weak edge regions. This branch effectively preserves texture details while suppressing noise responses. Compared to the gradient map, the phase map shows a more uniform overall intensity distribution and stronger structural continuity. The fused feature map visually integrates the advantages of both branches: it maintains clear edge contours while also capturing fine-grained texture structures. The target regions are more prominent in the fused result, whereas background noise is significantly suppressed. These results demonstrate that the AGPEO module can effectively combine gradient and phase information, enhancing edge clarity while preserving texture stability, thereby providing more discriminative and robust feature representations for infrared gas leak detection.
@@ 4.5.5 Multi-Scale Edge Representation Analysis Of MSEPM
To further investigate the edge representation capability of the proposed MSEPM, a four-scale visualization experiment was conducted, as illustrated in Figure. 11. The first scale (Scale 1) corresponds to the fused edge map generated by the AGPEO module without downsampling, preserving abundant fine-grained textures and weak gas boundaries. This layer effectively delineates faint and semi-transparent gas plumes against complex infrared backgrounds. After one downsampling operation (Scale 2), local noise and minor details are moderately suppressed, while continuous contour structures begin to emerge, reflecting a balance between detail retention and edge smoothness. The third scale (Scale 3) further aggregates regional responses, enhancing the integrity of the main gas contours and emphasizing medium-scale plume boundaries. At the deepest level (Scale 4), after three successive max-pooling operations, the representation becomes dominated by the global morphology of the gas plume, highlighting large-scale diffusion structures while discarding most fine textures. Overall, the visualization clearly demonstrates that shallow scales capture detailed boundary textures of weak gas plumes, whereas deeper scales progressively abstract these responses into structural and global representations. This hierarchical perception of edge information provides complementary guidance for both fine plume localization and global morphology interpretation in infrared gas leak detection.
@@ 4.5.7 Effect Of The Weighting Factor $\alpha$ In AGPEO
We conduct an ablation study on the fusion weight $\alpha$ within the AGPEO module of PEG-DRNet. Figure. 12. shows the detection performance under different settings of $\alpha$. When $\alpha$ is fixed, the performance varies noticeably across AP, AP50, and APSS, indicating that detection accuracy is sensitive to the choice of $\alpha$. Intermediate fixed values (e.g., $\alpha=0.4$$0.8$) improve small-object performance, while extreme settings exhibit inconsistent behavior across metrics. The gradient-only setting ($\alpha=1.0$) achieves relatively strong overall performance, whereas the phase-only setting ($\alpha=0$) yields lower accuracy, particularly for small objects. In contrast, the learnable $\alpha$ consistently outperforms all fixed configurations on AP, AP50{50}, and APS. This trend demonstrates that allowing $\alpha$ to be optimized during training leads to more stable and effective feature fusion than manually fixing its value.
In addition to the fixed-$\alpha$ experiments, we track the evolution of the learnable fusion weight $\alpha$ during the 200 training epochs, as shown in Figure 13. Initialized at 0.7, $\alpha$ rises steadily during the first  50 epochs and then gradually stabilizes within the range of approximately 0.77–0.79. The small fluctuations observed after convergence indicate that the network consistently identifies an effective weighting, avoiding extreme gradient-only or phase-only configurations. This adaptive behavior demonstrates the advantage of a learnable fusion weight, allowing AGPEO to balance multi-directional gradient and phase information across objects with diverse scales and edge characteristics.
Figure 13: Training evolution of the learnable fusion weight $\alpha$ in AGPEO. The curve shows that $\alpha$ rises during the first  50 epochs and stabilizes around 0.77–0.79, effectively balancing gradient and phase information.
@@ 5 Conclusion
This paper presents PEG-DRNet, a single-stage infrared gas leak detector that couples physics priors with content-adaptive feature learning. The Physics–Edge Hybrid Backbone incorporates a Gas Block that emulates diffusion–convection dynamics, and an AGPEO-based multi-scale edge perception module (MSEPM) that provides robust gradient–phase edge cues for semi-transparent, weak-contrast plumes. The proposed CASR-PAN further introduces content-adaptive sparse routing via AIMM-F and AIMM-S, selectively propagating informative cross-scale features while suppressing redundancy. Experiments on the IIG infrared gas dataset and LangGas dataset show that PEG-DRNet surpasses representative CNN- and Transformer-based detectors, achieving consistent improvements in $AP$, $AP_{50}$, and $AP_{S}$ over RT-DETR, while maintaining a lightweight design with 43.7 Gflops and 14.93M parameters, making it suitable for real-time deployment. Ablation studies confirm the complementary roles of physics priors, edge-guided enhancement, and adaptive routing. Future work will extend this framework toward temporal–spectral gas diffusion modeling, multi-gas joint detection, and edge-intelligent deployment in industrial monitoring scenarios.
@@ 6 CRediT Authorship Contribution Statement
Dongsheng Li: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Validation, Visualization, Writing – original draft, Writing – review & editing. Chaobo Chen: Funding acquisition, Supervision. Siling Wang: Software, Validation, Visualization, Writing – original draft. Song Gao: Funding acquisition, Supervision, Writing – review & editing.
@@ 8 Funding
This work was supported in part by the National Natural Science Foundation of China (No. 52572487), by the National Natural Science Foundation of China (No. 62303368), by the Qin Chuang Yuan “Scientist + Engineer” Team Construction Project (No. 2024QCY-KXJ-172).
@@ 9 List of Abbreviations
PANet path aggre-gation network
@@
IDAS id-aware scaling

### v2

@@ 4.5.2 Effective Receptive Field (ERF) Analysis Of Physics–Edge Hybrid Backbone
To further verify the physical interpretability and feature propagation capability of the proposed Physics–Edge Hybrid Backbone, we conducted an effective receptive field (ERF) visualization experiment to compare it with a conventional ResNet-based backbone under identical training and inference settings, as shown in Figure 7. Specifically, we employed the gradient-based ERF analysis method, where the gradient of the central activation with respect to the input pixels was accumulated and visualized as a heatmap (Ding et al., 2022). The visualization highlights high-response regions that contribute most significantly to the central neuron activation, thus reflecting the spatial extent and distribution of the model’s receptive field.
@@
Based on the quantitative statistics of the effective receptive field (ERF) distributions, a clear difference can be observed between the proposed Physics-Edge Hybrid Backbone and the conventional ResNet backbone. As shown in Figure 8, both the bar and line plots illustrate the relationship between the high-contribution area ratio and the activation threshold. Specifically, under low thresholds (e.g., 0.2 and 0.3), Physics-Edge Hybrid Backbone already achieves slightly higher area ratios than ResNet, indicating a broader response range even in the early activation stage. As the threshold increases to 0.5 and 0.99, this gap becomes more pronounced — the high-contribution area ratio of Gas Block reaches 0.047 and 0.340, compared with 0.036 and 0.255 for ResNet, respectively. This trend reveals that the proposed Gas Block possesses a larger and more spatially continuous receptive field, allowing information to propagate more effectively across distant regions. Consequently, the network is better able to integrate global contextual cues while preserving fine structural details, consistent with the convection–diffusion-Inspired design of Physics-Edge Hybrid Backbone.
@@ 4.5.3 Backbone-Level Comparison With Representative Architectures
To comprehensively evaluate the effectiveness of the proposed Physics-Edge Hybrid Backbone, we conducted a comparative study against several representative models, including EfficientNet (Tan and Le, 2019), MambaOut (Yu and Wang, 2025), RepViT (Wang et al., 2024), CSWin Transformer (Dong et al., 2022),ConvNeXtV2 (Woo et al., 2023), RMT (Fan et al., 2024), EfficientFormerV2 (Li et al., 2023), LSKNet (Li et al., 2023), and VanillaNet (Chen et al., 2023), highlighting the performance improvements achieved by our method, as shown in Figure 9.
@@ 4.5.4 Visualization Analysis Of AGPEO Edge Representation
To validate the feature extraction capability of the AGPEO module, we present a comparative visualization of the original infrared image, the gradient feature map, the phase feature map, and the fused result, as shown in Figure 10. It can be observed that the gradient branch, employing multi-directional Sobel operators, effectively highlights prominent edge responses, emphasizing both the gas leak regions and the structural contours of the equipment. However, due to the presence of temperature gradients and noise in the infrared image background, the gradient feature map also exhibits noticeable spurious edges in some background areas, resulting in partially fragmented boundaries. The phase branch utilizes Gabor convolutional kernels to perform local phase modulation, producing phase feature maps that are smoother in weak edge regions. This branch effectively preserves texture details while suppressing noise responses. Compared to the gradient map, the phase map shows a more uniform overall intensity distribution and stronger structural continuity. The fused feature map visually integrates the advantages of both branches: it maintains clear edge contours while also capturing fine-grained texture structures. The target regions are more prominent in the fused result, whereas background noise is significantly suppressed. These results demonstrate that the AGPEO module can effectively combine gradient and phase information, enhancing edge clarity while preserving texture stability, thereby providing more discriminative and robust feature representations for infrared gas leak detection.
@@ 4.5.5 Multi-Scale Edge Representation Analysis Of MSEPM
To further investigate the edge representation capability of the proposed MSEPM, a four-scale visualization experiment was conducted, as illustrated in Figure 11. The first scale (Scale 1) corresponds to the fused edge map generated by the AGPEO module without downsampling, preserving abundant fine-grained textures and weak gas boundaries. This layer effectively delineates faint and semi-transparent gas plumes against complex infrared backgrounds. After one downsampling operation (Scale 2), local noise and minor details are moderately suppressed, while continuous contour structures begin to emerge, reflecting a balance between detail retention and edge smoothness. The third scale (Scale 3) further aggregates regional responses, enhancing the integrity of the main gas contours and emphasizing medium-scale plume boundaries. At the deepest level (Scale 4), after three successive max-pooling operations, the representation becomes dominated by the global morphology of the gas plume, highlighting large-scale diffusion structures while discarding most fine textures. Overall, the visualization clearly demonstrates that shallow scales capture detailed boundary textures of weak gas plumes, whereas deeper scales progressively abstract these responses into structural and global representations. This hierarchical perception of edge information provides complementary guidance for both fine plume localization and global morphology interpretation in infrared gas leak detection.
@@ 4.5.7 Effect of the Weighting Factor $\alpha$ in AGPEO
We conduct an ablation study on the fusion weight $\alpha$ within the AGPEO module of PEG-DRNet. Figure 12. shows the detection performance under different settings of $\alpha$. When $\alpha$ is fixed, the performance varies noticeably across AP, AP50, and APSS, indicating that detection accuracy is sensitive to the choice of $\alpha$. Intermediate fixed values (e.g., $\alpha=0.4$$0.8$) improve small-object performance, while extreme settings exhibit inconsistent behavior across metrics. The gradient-only setting ($\alpha=1.0$) achieves relatively strong overall performance, whereas the phase-only setting ($\alpha=0$) yields lower accuracy, particularly for small objects. In contrast, the learnable $\alpha$ consistently outperforms all fixed configurations on AP, AP50{50}, and APS. This trend demonstrates that allowing $\alpha$ to be optimized during training leads to more stable and effective feature fusion than manually fixing its value.
In addition to the fixed-$\alpha$ experiments, we track the evolution of the learnable fusion weight $\alpha$ during the 200 training epochs, as shown in Figure 13. Initialized at 0.7, $\alpha$ increases during the first $\sim$50 epochs and then enters a stable regime, where it remains within a narrow range of approximately 0.77–0.79 for the remainder of training. The magnitude of variation after this point is small and does not exhibit a monotonic increasing or decreasing trend, indicating practical convergence of the fusion weight. This behavior suggests that the network consistently identifies an effective balance between gradient and phase cues, rather than drifting toward extreme gradient-only or phase-only configurations. Overall, the observed stability supports the use of a learnable fusion weight for adaptive and robust edge representation.
Figure 13: Training evolution of the learnable fusion weight $\alpha$ in AGPEO during the 200 epochs. After an initial increasing phase, $\alpha$ fluctuates within a narrow range, suggesting a stable fusion preference during training.
@@ 5 Conclusion
This paper presents PEG-DRNet, a single-stage infrared gas leak detector that couples physics priors with content-adaptive feature learning. The Physics–Edge Hybrid Backbone incorporates a Gas Block that emulates diffusion–convection dynamics, and an AGPEO-based multi-scale edge perception module (MSEPM) that provides robust gradient–phase edge cues for semi-transparent, weak-contrast plumes. The proposed CASR-PAN further introduces content-adaptive sparse routing via AIMM-F and AIMM-S, selectively propagating informative cross-scale features while suppressing redundancy. Experiments on the IIG infrared gas dataset and LangGas dataset show that PEG-DRNet surpasses representative CNN- and Transformer-based detectors, achieving consistent improvements in $AP$, $AP_{50}$, and $AP_{S}$ over RT-DETR, while maintaining a lightweight design with 43.7 Gflops and 14.93M parameters, making it suitable for real-time deployment. Ablation studies confirm the complementary roles of physics priors, edge-guided enhancement, and adaptive routing. Future work will extend this framework toward temporal–spectral gas diffusion modeling, multi-gas joint detection, and edge-intelligent deployment in industrial monitoring scenarios.In addition, recent advances in vision foundation models and large-scale multimodal models have demonstrated remarkable generalization capabilities under limited or even zero-shot supervision. Exploring how such foundation models can be effectively adapted or distilled for infrared gas leak detection—particularly to alleviate data scarcity and improve robustness across unseen environments—represents a promising direction for future research.
@@ 6 CRediT Authorship Contribution Statement
Dongsheng Li: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Validation, Visualization, Writing – original draft, Writing – review & editing. Tianli Ma: Funding acquisition, Supervision. Siling Wang: Software, Validation, Visualization, Writing – original draft. Beibei Duan:Validation, Supervision. Song Gao: Funding acquisition, Supervision, Writing – review & editing.
@@ 8 Funding
This work was supported in part by the National Natural Science Foundation of China (No. 62303368), by the National Foreign Experts Project (No. H20251091), by the Qin Chuang Yuan “Scientist + Engineer” Team Construction Project (No. 2024QCY-KXJ-172).
@@ 9 List of Abbreviations
Table 10: List of abbreviations used in this paper
@@
PANet path aggregation network
@@
IDAS identity-aware scaling